# Foldings in relatively hyperbolic groups

### v1

@@ 2.2 $(G,\mathbb{P})$-carrier graphs of groups
(2) In this case $w:=\alpha(e)\in VC_{i}$ for some non-trivial star $C_{i}$, $A_{e}=\alpha_{e}(A_{e})\in P_{m_{i}}$ and $\omega_{e}(A_{e})=g_{e}^{-1}\alpha_{e}(A_{e})g_{e}\leq g_{e}^{-1}P_{m_{i}}g_{e}$. By assumption $g_{e}\alpha_{e}(A_{e})g_{e}^{-1}$ is a proper subgroup of some parabolic subgroup of $A_{v}$, i.e. of some subgroup $A_{v}\cap g_{e}^{-1}P_{m_{i}}g_{e}$. Replace $A_{e}$ by $A_{e}^{\prime}:=g_{e}(A_{v}\cap g_{e}^{-1}P_{m_{i}}g_{e})g_{e}^{-1}=g_{e}A_{v}g_{e}^{-1}\cap P_{m_{i}}$, $A_{w}$ by $\langle A_{w},A_{e}^{\prime}\rangle$ (and adjust the peripheral star $C_{i}$ as discussed above). Note that on the graph of group level this corresponds to a fold of type IIA in the sense of [8].
@@
If $C_{i}=C_{j}$ (and possibly $w_{1}=w_{2}$) then there exists a path $(f_{1},\ldots,f_{k})$ of length at most $2$ in $C_{i}$ from $w_{2}$ to $w_{1}$. Let $g^{\prime}=g_{f_{1}}\cdot\ldots\cdot g_{f_{k}}\in P_{m_{i}}$. Clearly $\bar{g}:=g_{e_{1}}gg_{e_{2}}^{-1}g^{\prime}\in P_{m_{i}}$ is an element represented by a closed $\mathbb{A}$-path at $w_{1}$. We remove the edge $e_{2}$ and replace $A_{w_{1}}$ by $\langle A_{w_{1}},\bar{g}\rangle$ (and adjust the peripheral star $C_{i}$ as discussed above), see Figure 3. Subsequently we remove $w_{2}$ and its only adjacent edge if $w_{2}$ is of valence $1$.
@@
(1) In the last step the elements of essential edges are modified to guarantee they are of minimal possible length, i.e. that 4. of Definition 2.9 holds. By Lemma 1.10 this also implies that (1) holds.
Suppose that $\lvert pg_{e}\rvert_{X}<\lvert g_{e}\rvert_{X}$ for some essential edge $e\in EA$ with $w=\alpha(e)\in VC_{i}$ and $p\in P_{m_{i}}$. Let $g$ be the element represented by the unique edge in $C_{i}$ from $c_{i}$ to $w$. Remove the edge $e$ and introduce a new vertex $w^{\prime}$ and new edges $f^{\prime}$ and $e^{\prime}$ with $\alpha(f^{\prime})=c_{i}$, $\omega(f^{\prime})=\alpha(e^{\prime})$, $\omega(e^{\prime})=v$, $g_{f^{\prime}}=gp^{-1}$ and $g_{e^{\prime}}=pg_{e}$, see Figure 6.
@@
    The number of free factors and the free rank do not increase.
@@ 2.3 Metrics on fundamental groups of $(G,\mathbb{P})$-carrier graphs
Let $t=(a_{0},e_{1},a_{1},\ldots,e_{k},a_{k})$ be an $\mathbb{A}$-path, $v_{0}:=\alpha(e_{1})$, and further $v_{m}:=\omega(e_{m})$ for $m\in\{1,\ldots,k\}$. Let $s=s_{0}\ldots s_{2k}$ be a path in $\operatorname{Cay}(G,X\cup\mathcal{P})$ such that the following hold:
@@
If $s_{1}$ is itself the realization of some $\mathbb{A}$-subpath $t^{\prime}$ of $t$, there is nothing to show. Hence, one of the following cases occurs:
@@
    If on the other hand $e$ is essential, let $s^{\prime}:=r_{1}$. $s^{\prime}$ is a realization of some $\mathbb{A}$-subpath of $t$ and
@@
In fact, it is in general not possible to give any upper bound on $\lvert s\rvert_{X}^{\mathcal{A}}$ in terms of $\lvert t\rvert_{X}^{\mathcal{A}}$ if $s$ is not required to be a full $\mathbb{A}$-subpath of $t$.
@@
Let $T$ be a spanning tree of $A$ that contains the stars underlying all non-trivial $\mathcal{C}_{i}$. For each $v\in VA$ let $\gamma_{v}$ the unique reduced $\mathbb{A}$-path with trivial vertex elements from $v_{0}$ to $v$ in $T$. Let $\mathbb{O}_{\mathcal{A}}$ be the set of non-trivial subgroups $[\gamma_{c_{i}}A_{c_{i}}\gamma_{c_{i}}^{-1}]$ of $\pi_{1}(\mathbb{A},v_{0})$ with $i\in\{1,\ldots,k\}$.
@@ 2.4 $M$-normal $(G,\mathbb{P})$-carrier graphs of groups
We will be exclusively interested in realizations of $\mathbb{A}$-paths of $M$-prenormal $(G,\mathbb{P})$-carrier graphs, for such a realization. Throughout this section $(G,\mathbb{P})$ is a torsion-free relatively hyperbolic group with symmetric finite generating set $X$, $(\mathcal{A},((\mathcal{C}_{i},c_{i}))_{1\leq i\leq k})$ is an $M$-prenormal $(G,\mathbb{P})$-carrier graph of groups and $\nu:=\nu(G,\mathbb{P},X,M)$ is the constant from Lemma & Definition 1.10.
@@
    $e$ is an edge of some $s_{2i}$ (coming from an element of an essential vertex group) with $\operatorname{len}_{X}(e)\geq\nu/2$.
@@
    $e$ is the single edge of some $s_{2i}$ (coming from a non-trivial element of an essential vertex group, possibly with $\operatorname{len}_{X}(e)<\nu/2$).
@@
$\operatorname{len}_{X}(p_{1}),\operatorname{len}_{X}(p_{2})\leq\nu.$
Let $\mathcal{A}$ be an $M$-prenormal $(G,\mathbb{P})$ carrier graph of groups. $\mathcal{A}$ is called $M$-normal if every $\mathbb{A}$-path is equivalent to an $\mathbb{A}$-path of minimal $X\cup\mathcal{P}$-length that has distinguished $\mathcal{P}$-components.
@@
Suppose $s=s_{1}ps_{2}$ where $p$ is a distinguished edge of $s$ of the second type, i.e a $P_{m}$-edge with $\operatorname{len}(p)_{X}\geq\nu/2$.
@@
    $p^{\prime}$ is a $P_{m}$-edge connected to $p$ with margin: $p_{-}$ instead of $\alpha(p)$ (etc.)?
@@
Note that condition 4 of Definition 2.9 implies that if $A_{v}\cap P_{i}\neq 1$ for some $i$ then there exists $e\in E_{v}$ with $\omega_{e}(A_{e})=A_{v}\cap P_{i}$ and $g_{e}=1$. It is moreover a consequence of Lemma 2.21 (See Remark 1.11) that this also holds if there exists $a\in A_{v}$ such that a geodesic representing $a$ starts or ends with an element of $P_{i}$ of $X$-length at least $\nu/2$.
Lemma 2.21 gives some justification for the definition of a distinguished edge in Definition 2.19. Indeed, let $t=(\ldots,a,\ldots)$ be an $\mathbb{A}$-path where $a$ is an element of some essential vertex group, $s$ is a realization of $t$ and $p$ is a distinguished edge in the geodesic subpath of $s$ associated with $a$. Then it follows from the previous lemma, that $t$ can be modified into an $\mathbb{A}$-path $t^{\prime}$ with $\nu_{\mathcal{A}}(t^{\prime})=\nu_{\mathcal{A}}(t)$ by replacing the $\mathbb{A}$-subpath $(a)$ of $t$ with some $\mathbb{A}$-path $(a_{1},e,a^{\prime},e^{-1},a_{2})$ for an essential edge $e$, such that for some realization $s^{\prime}$ of $t^{\prime}$ with the same endpoints as $s$, $p$ is connected to the distinguished edge $p^{\prime}$ of $s^{\prime}$ corresponding to $a^{\prime}$ and the $X$-distance of the endpoints of $p$ and $p^{\prime}$ is at most $3\nu$.
@@
Assume $\mathcal{A}$ is not $M$-normal and show that this implies that 1.-4 holds.
@@
We can therefore assume that all reduced paths of non-peripheral edges labeled by elements of some $P_{m}$ are simple (and not closed). We may further assume that no trivial edge is labeled by a non-trivial element of some $P_{m}$ as otherwise 3. occurs.
Since $\mathcal{A}$ is not $M$-normal, there is some $\mathbb{A}$-path $t$ of minimal $X\cup\mathcal{P}$-length which is not equivalent to an $\mathbb{A}$-path of minimal $X\cup\mathcal{P}$-length that has distinguished $\mathcal{P}$-components. Let $s$ be a realization of $t$ and suppose that $s$ has the minimal number of distinguished edges of all realizations of $\mathbb{A}$-paths of minimal $X\cup\mathcal{P}$-length equivalent to $t$. By assumption there is a $P_{m}$-component $p$ of $s$ for some $m$ with $\operatorname{len}_{X}(p)\geq 3\nu$ which is not of the form $p_{1}ep_{2}$ for some distinguished edge $e$ of $s$ with $\operatorname{len}_{X}(p_{1}),\operatorname{len}_{X}(p_{2})\leq\nu$.
@@
    $p$ is of the form $p_{1}ep_{2}$, where $e$ is a distinguished edge of $s$, $p_{1}$ and $p_{2}$ have no distinguished edges, and $\operatorname{len}_{X}(p_{1})>\nu$ or $\operatorname{len}_{X}(p_{2})>\nu$.
We may assume that 3. does not occur, in particular no trivial edge is labeled by a non-trivial peripheral element.
@@
  • $p_{1}$ ($p_{k}$) is a suffix (prefix) of length at most $\nu/2$ of some geodesic representing an element of an essential vertex or the element of some essential edge.
  • $p_{i}$ is a geodesic representing the element of some essential edge (and is therefore of length at most $\nu/2$) for all $i\in\{2,\ldots,k-1\}$.
As $\operatorname{len}_{X}(p)\geq 3\nu$ it follows that $k\geq 6$. Thus, for some $i$ the subpath $p_{i}p_{i+1}$ is a subpath of $X$-length at most $\nu$ that corresponds to a sub-$\mathbb{A}$-path of $t$ connecting two essential vertices which is simple by the above remark. Hence, 1. occurs.
@@
If $\hat{t}$ ist the $\mathbb{A}$-subpath of $t$ representing $q$ then $t^{\prime}=\hat{t}\cdot(1,e,1)$ is an $\mathbb{A}$ path connecting the $P_{m}$ star corresponding to $e_{1}$ to the $P_{m}$-star $P_{i}$. Note that $q$ is a realization of $t^{\prime}$ as $g_{e}=1$. Note further that $\hat{t}$ and therefore also $t^{\prime}$ does not contain a peripheral edge. There are three cases to consider:
  • $t^{\prime}$ is reduced and non-degenerate. Then the same argument as in the first case shows that $\mathcal{A}$ fulfills 1. or 4.
  • $t^{\prime}$ is not reduced as $\hat{t}$ ist of the form $(\ldots,e^{-1},1)$ and $t^{\prime}$ reduces to a non-degenerate path. Then the argument from the first case applied to the obtained path shows that $\mathcal{A}$ fulfills 1. or 4.
  • $t^{\prime}$ is not reduced as $\hat{t}$ ist of the form $(\ldots,e^{-1},1)$ and $t^{\prime}$ reduces to a degenerate path $(1)$. using Lemma 2.21, it is now possible to construct a realization $s^{\prime}$ of some $\mathbb{A}$-path which is equivalent to $t$ and has the same $X\cup\mathcal{P}$-length as $t$, such that $s^{\prime}$ has fewer peripheral edges than $s$, a contradiction to the choice of $s$. This is achieved by pushing part of $e_{2}$ past $e^{-1}$ which is possible as $g_{e}=1$.
(3) Lastly, suppose $p$ is of the form $p_{1}ep_{2}$, where $e$ is a distinguished edge of $s$, $p_{1}$ and $p_{2}$ have no distinguished edges, and w.l.o.g. $\operatorname{len}_{X}(p_{2})>\nu$. Applying Lemma 2.21, it can be assumed that $e$ corresponds to a maximal $\mathbb{C}_{i}$-subpath of $t$. Indeed, if $e$ is a distinguished edge of the second or third type, i.e. labeled by an element of $P_{m}$ that is the terminal edge of a geodesic representing an element $a$ of some essential vertex group, then the element of the essential edge $e$ connecting $v$ and the appropriate carrier star $\mathcal{C}_{i}$ must be trivial, see Remark 2.22. It follows that $p_{2}$ is an initial path of a realization of some $\mathbb{A}$-path $t^{\prime}$ with $t^{\prime}_{-}\in VC_{i}$. As in (1) we may assume that $t^{\prime}$ has only essential edges, free edges with trivial edge element and possibly some non-trivial vertex group element $a$ of an essential vertices at the very end such that $p_{2}$ reads the initial edge of the geodesic representing $a$ which is of $X$-length at most $\nu/2$.
The same argument as in case (1) shows that 1. occurs. $\square$

### v2

@@ 2.2 $(G,\mathbb{P})$-carrier graphs of groups
(2) In this case $w:=\alpha(e)\in VC_{i}$ for some non-trivial star $C_{i}$, $A_{e}=\alpha_{e}(A_{e})\in P_{m_{i}}$ and $\omega_{e}(A_{e})=g_{e}^{-1}\alpha_{e}(A_{e})g_{e}\leq g_{e}^{-1}P_{m_{i}}g_{e}$. By assumption $g_{e}^{-1}\alpha_{e}(A_{e})g_{e}$ is a proper subgroup of some parabolic subgroup of $A_{v}$, i.e. of some subgroup $A_{v}\cap g_{e}^{-1}P_{m_{i}}g_{e}$. Replace $A_{e}$ by $A_{e}^{\prime}:=g_{e}(A_{v}\cap g_{e}^{-1}P_{m_{i}}g_{e})g_{e}^{-1}=g_{e}A_{v}g_{e}^{-1}\cap P_{m_{i}}$, $A_{w}$ by $\langle A_{w},A_{e}^{\prime}\rangle$ (and adjust the peripheral star $C_{i}$ as discussed above). Note that on the graph of group level this corresponds to a fold of type IIA in the sense of [9].
@@
Suppose that $C_{i}=C_{j}$. If $w_{1}\neq w_{2}$ then there exists a path $(f_{1},f_{2})$ in $C_{i}$ from $w_{2}$ to $w_{1}$ and we set $g^{\prime}=g_{f_{1}}\cdot g_{f_{k}}\in P_{m_{i}}$. If $w_{1}=w_{2}$ we set $g^{\prime}=1$. In both cases $\bar{g}:=g_{e_{1}}gg_{e_{2}}^{-1}g^{\prime}\in P_{m_{i}}$ is an element represented by a closed $\mathbb{A}$-path at $w_{1}$. We remove the edge $e_{2}$ and replace $A_{w_{1}}$ by $\langle A_{w_{1}},\bar{g}\rangle$ (and adjust the peripheral star $C_{i}$ as discussed above), see Figure 3. Subsequently we remove $w_{2}$ and its only adjacent edge if $w_{2}$ is of valence $1$.
@@
(1) In this step the elements of essential edges are modified to guarantee they are of minimal possible length, i.e. that 4. of Definition 2.9 holds. By Lemma 1.10 this also implies that (1) holds.
Suppose that $\lvert pg_{e}a\rvert_{X}<\lvert g_{e}\rvert_{X}$ where $e\in EA$ is some essential edge with $w=\alpha(e)\in VC_{i}$, $p\in P_{m_{i}}$ and $a\in A_{v}$. Let $g$ be the element represented by the unique edge in $C_{i}$ from $c_{i}$ to $w$. Remove the edge $e$ and introduce a new vertex $w^{\prime}$ and new edges $f^{\prime}$ and $e^{\prime}$ with $\alpha(f^{\prime})=c_{i}$, $\omega(f^{\prime})=\alpha(e^{\prime})$, $\omega(e^{\prime})=v$, $g_{f^{\prime}}=gp^{-1}$ and $g_{e^{\prime}}=pg_{e}a$, see Figure 6.
@@
In a final step one guarantees that 7. of Definition 2.6 is guaranteed by passing to the appropriate sub-$\mathbb{A}$-graph. There are some choices in this process and an inspection of all cases shows that the process can performed such that the vertex group of the base vertex is not conjugated. In this case, by choosing the base vertex to be image of the original base vertex under the above operations, the resulting $(G,\mathbb{P})$-carrier graph represents the same subgroup.
@@
    The number of free factors and the relative rank do not increase.
@@ 2.3 Metrics on fundamental groups of $(G,\mathbb{P})$-carrier graphs
Let $t=(a_{0},e_{1},a_{1},\ldots,e_{k},a_{k})$ be an $\mathbb{A}$-path, $v_{0}:=\alpha(e_{1})$ and $v_{m}:=\omega(e_{m})$ for $m\in\{1,\ldots,k\}$. Let $s=s_{0}\ldots s_{2k}$ be a path in $\operatorname{Cay}(G,X\cup\mathcal{P})$ such that the following hold:
@@
If $s_{1}$ is itself the realization of some $\mathbb{A}$-subpath $t^{\prime}$ of $t$, there is nothing to show. Hence, one of the following two cases occurs:
@@
    If on the other hand $e$ is essential, let $^{\prime}:=r_{1}. ss^{\prime}$ is a realization of some $\mathbb{A}$-subpath of $t$ and
@@
Indeed, for an $\mathbb{A}$-path $t$ consisting of a single $\mathbb{C}_{i}$-path $t=(1,e_{1},1,e_{2},1)$ with $\mathbb{C}_{i}$-subpath $s=(1,e_{1},1)$ it is in general not possible to give any upper bound on $\lvert s\rvert_{X}^{\mathcal{A}}=|g_{e_{1}}|_{X}$ in terms of $\lvert t\rvert_{X}^{\mathcal{A}}=|g_{e_{1}}g_{e_{2}}|_{X}$.
@@
Let $T$ be a spanning tree of $A$ that contains the stars underlying all non-trivial $\mathcal{C}_{i}$. For each $v\in VA$ let $\gamma_{v}$ the unique reduced $\mathbb{A}$-path with trivial vertex elements from $v_{0}$ to $v$ in $T$. Let $\mathbb{O}_{\mathcal{A}}$ be the set of non-trivial subgroups $\gamma_{c_{i}}A_{c_{i}}\gamma_{c_{i}}^{-1}]$ of $\pi_{1}(\mathbb{A},v_{0})$ with $i\in\{1,\ldots,k\}$.
@@ 2.4 $M$-normal $(G,\mathbb{P})$-carrier graphs of groups
Throughout this section $(G,\mathbb{P})$ is a torsion-free relatively hyperbolic group with symmetric finite generating set $X$, $(\mathcal{A},((\mathcal{C}_{i},c_{i}))_{1\leq i\leq k})$ is an $M$-prenormal $(G,\mathbb{P})$-carrier graph of groups and $\nu:=\nu(G,\mathbb{P},X,M)$ is the constant from Lemma & Definition 1.10.
@@
    $e$ is an edge of some $s_{2i}$ (coming from an element of an essential vertex group) with $\operatorname{len}_{X}(e)>\frac{3}{2}\nu$.
@@
    $e$ is the single edge of some $s_{2i}$ (coming from a non-trivial element of an essential vertex group, possibly with $\operatorname{len}_{X}(e)\leq\frac{3}{2}\nu$).
@@
Definition 2.20.
Let $t=(a_{0},e_{1},a_{1},\ldots,e_{k},a_{k})$ be a reduced $\mathbb{A}$-path. $t$ is called tame if the following hold:
  1. 1.
    Suppose that $e_{i}$ and $\omega(e_{i})$ are essential, that $\alpha(e_{i})\in VC_{j}$ where $\mathcal{C}_{j}$ is a $P_{m_{j}}$-carrier star and $s=s_{0}s_{1}s_{2}$ is a realization of the $\mathbb{A}$-path $(p,e_{i},a_{i})$ with $p\in P_{m_{i}}\neq 1$. Then no two distinguised edges of $s$ are connected.
  2. 2.
    Suppose that $e_{i}$ and $\alpha(e_{i})$ are essential, that $\omega(e_{i})\in VC_{j}$ where $\mathcal{C}_{j}$ is a $P_{m_{j}}$-carrier star and $s=s_{0}s_{1}s_{2}$ is a realization of the $\mathbb{A}$-path $(a_{i-1},e_{i},p)$ with $p\in P_{m_{i}}\neq 1$. Then no two distinguised edges of $s$ are connected.
Note that condition 1. of Definition 2.20 is satisfied for $t$ iff condition 2. is statisfied for $t^{-1}$, in particular $t$ is tame iff $t^{-1}$ is tame. The following lemma implies that we can always assume that elements are represented by tame $\mathbb{A}$-paths.
Lemma 2.21.
Any reduced $\mathbb{A}$-path is equivalent to a tame reduced $\mathbb{A}$-path.
Proof.
Suppose that a reduced $\mathbb{A}$-path $t=(a_{0},e_{1},a_{1},\ldots,e_{k},a_{k})$ is not tame. Possibly after replacing $t$ by $t^{-1}$ we may assume that there exists $i$ such that $e_{i}$ and $\omega(e_{i})$ are essential, that $\alpha(e_{i})\in VC_{j}$ where $\mathcal{C}_{j}$ is a $P_{m_{j}}$-carrier star and that some realization $s=s_{0}s_{1}s_{2}$ of some $\mathbb{A}$-path $(p,e_{i},a_{i})$ with $p\in P_{m_{i}}\neq 1$ has connected distinguished components.
If follows from Corollary 1.12 that there exists some
$a\in A_{\omega(e_{i})}\cap P_{m_{j}}^{g_{j}}=\hbox{Im}(\omega_{e_{i}})=g_{j}^{-1}A_{e_{i}}g_{j}$
such that $|aa_{i}|_{X}<|a_{i}|_{X}$ and $|aa_{i}|_{X\cup\mathcal{P}}\leq|a_{i}|_{X\cup\mathcal{P}}$. Write $a=g_{j}^{-1}a^{\prime}g_{j}=\omega_{e_{i}}(a^{\prime})$ with $a^{\prime}\in A_{e_{i}}\subset P_{m_{j}}$ and in $t$ replace $a_{i}$ by $aa_{i}=\omega_{e_{i}}(a^{\prime})a_{i}$ and $a_{i-1}$ by $a_{i-1}\alpha_{e_{i}}(a^{\prime})=a_{i-1}a^{\prime}$. This clearly yields an equivalent $\mathbb{A}$-path. Perform this operation as often as possible, the result is a tame $\mathbb{A}$-path equivalent to the original $\mathbb{A}$-path $t$. $\square$
Remark 2.22.
It is in general not possible to find a tame $\mathbb{A}$-path that is of minimal $(X\cup\mathcal{P})$-length in its equivalence class. Indeed, in the construction of the tame $\mathbb{A}$-graph a trivial element picked up in a peripheral star might be replaced by a non-trivial element.
@@
$\operatorname{len}_{X}(p_{1}),\operatorname{len}_{X}(p_{2})\leq 5\nu.$
Let $\mathcal{A}$ be an $M$-prenormal $(G,\mathbb{P})$ carrier graph of groups. $\mathcal{A}$ is called $M$-normal if every tame $\mathbb{A}$-path has distinguished $\mathcal{P}$-components.
Remark 2.24.
In Definition 2.23 one could choose $p$ to be an arbitrary $P_{i}$-path of $s$ as any such path is a $\mathcal{P}$-component of the realization of some tame $\mathbb{A}$-path
@@
Suppose $s=s_{1}ps_{2}$ where $p$ is a distinguished edge of $s$ of the second type, i.e a $P_{m}$-edge with $\operatorname{len}(p)_{X}\geq\frac{3}{2}\nu$.
@@
    $p^{\prime}$ is a $P_{m}$-edge connected to $p$ with
@@
$\leq\lvert d_{X}((s_{1}^{\prime})_{-},(s_{1}^{\prime})_{+})-d_{X}((s_{1})_{-},(s_{1})_{+})\rvert+\lvert d_{X}((r_{1})_{-},(r_{1})_{+})\rvert$
@@
Note that condition 4 of Definition 2.9 implies that if $A_{v}\cap P_{i}\neq 1$ for some $i$ then there exists $e\in E_{v}$ with $\omega_{e}(A_{e})=A_{v}\cap P_{i}$ and $g_{e}=1$. It is moreover a consequence of Lemma 2.25 that this also holds if there exists $a\in A_{v}$ such that a geodesic word representing $a$ starts or ends with element of $P_{i}$ of $X$-length at least $\nu/2$.
Lemma 2.25 gives some justification for the definition of a distinguished edge in Definition 2.19. Indeed, let $t=(\ldots,a,\ldots)$ be an $\mathbb{A}$-path where $a$ is an element of some essential vertex group, $s$ is a realization of $t$ and $p$ is a distinguished edge in the geodesic subpath of $s$ associated with $a$. Then it follows from Lemma 2.25, that $t$ can be modified into an $\mathbb{A}$-path $t^{\prime}$ with $\nu_{\mathcal{A}}(t^{\prime})=\nu_{\mathcal{A}}(t)$ by replacing the $\mathbb{A}$-subpath $(a)$ of $t$ with some $\mathbb{A}$-path $(a_{1},e,a^{\prime},e^{-1},a_{2})$ for an essential edge $e$, such that for some realization $s^{\prime}$ of $t^{\prime}$ with the same endpoints as $s$, $p$ is connected to the distinguished edge $p^{\prime}$ of $s^{\prime}$ corresponding to $a^{\prime}$ and the $X$-distance of the endpoints of $p$ and $p^{\prime}$ is at most $2\nu$.
@@
Assume $\mathcal{A}$ is not $M$-normal and show that this implies that one of 1.-4. holds.
@@
We can therefore assume that all reduced paths of non-peripheral edges labeled by elements of some $P_{m}$ are simple (and not closed).
Since $\mathcal{A}$ is not $M$-normal, there is some tame $\mathbb{A}$-path $t$ which does not have distinguished $\mathcal{P}$-components. Let $s$ be a realization of $t$. By assumption there is a $P_{m}$-component $p$ of $s$ for some $m$ with $\operatorname{len}_{X}(p)\geq 10\nu$ which is not of the form $p_{1}ep_{2}$ for some distinguished edge $e$ of $s$ with $\operatorname{len}_{X}(p_{1}),\operatorname{len}_{X}(p_{2})\leq 5\nu$.
@@
    $p$ is of the form $p_{1}ep_{2}$, where $e$ is a distinguished edge of $s$, $p_{1}$ and $p_{2}$ have no distinguished edges, and $\operatorname{len}_{X}(p_{1})>5\nu$ or $\operatorname{len}_{X}(p_{2})>5\nu$.
We may assume that 3. of Lemma 2.27 does not occur, in particular no free edge is labeled by a non-trivial peripheral element.
@@
  • $p_{1}$ ($p_{k}$) is a suffix (prefix) of length at most $\frac{3}{2}\nu$ of some geodesic representing an element of an essential vertex or of the element of some essential edge.
  • For all $i\in\{2,\ldots,k-1\}$ the path $p_{i}$ is a geodesic representing the element of some essential edge (and is therefore of length at most $\nu/2$).
As $\operatorname{len}_{X}(p)\geq 10\nu\geq 5\nu$ it follows that $k\geq 6$. Thus, for some $i$ the subpath $p_{i}p_{i+1}$ is a subpath of $X$-length at most $\nu$ that corresponds to a sub-$\mathbb{A}$-path of $t$ connecting two essential vertices which is simple by the discussion at the beginning of the proof. Hence, 1. occurs.
@@
If $t^{\prime}$ ist the $\mathbb{A}$-subpath of $t$ representing $q$ then $\hat{t}=t^{\prime}\cdot(1,e,1)$ is an $\mathbb{A}$ path connecting the $P_{m}$ star corresponding to $e_{1}$ to the $P_{m}$-star $C_{i}$. Note that $q$ is a realization of $t^{\prime}$ as $g_{e}=1$. Note further that $t^{\prime}$ and therefore also $\hat{t}$ does not contain a peripheral edge.
Note that the tameness of $t$ imply that $\hat{t}$ is reduced. Thus $\hat{t}$ is reduced and non-degenerate and the same argument as in the first case shows that $\mathcal{A}$ fulfills 1. or 4.
(3) Lastly, suppose $p$ is of the form $p_{1}ep_{2}$, where $e$ is a distinguished edge of $s$, $p_{1}$ and $p_{2}$ have no distinguished edges, and w.l.o.g. $\operatorname{len}_{X}(p_{2})>5\nu$. We can now argue as in (1) with $p_{2}$ playing the role of $p$. $\square$